d^2y/-6 dy/dx+9y=1+x+x^2
1
Expert's answer
2021-09-17T03:45:44-0400

The homogeneous differential equation


"\dfrac{d^2y}{dx^2}-6\dfrac{dx}{dy}+9y=0"

The corresponding (auxiliary) equation


"r^2-6r+9=0"

"(r-3)^2=0"

"r_1=r_2=3"

The general solution of the homogeneous differential equation


"y_h=c_1e^{3x}+c_2xe^{3x}"

Find the particular solution of the nonhomogeneous differential equation


"y_p=A+Bx+Cx^2"

"\dfrac{dy_p}{dx}=B+2Cx"

"\dfrac{d^2y_p}{dx^2}=2C"

Substitute


"2C-6(B+2Cx)+9(A+Bx+Cx^2)=1+x+x^2"

"9C=1""-12C+9B=1""2C-6B+9A=1"

"A=\dfrac{7}{27}, B=\dfrac{7}{27}, C=\dfrac{1}{9}"

"y_p=\dfrac{7}{27}+\dfrac{7}{27}x+\dfrac{1}{9}x^2"

The general solution of the non homogeneous differential equation is

"y=y_h+y_p"

The solution of the given differential equation is


"y=c_1e^{3x}+c_2xe^{3x}+\dfrac{7}{27}+\dfrac{7}{27}x+\dfrac{1}{9}x^2"